OpenStudy (anonymous):

What is the area of a circle that has the radius 5?

OpenStudy (anonymous):

78.53

OpenStudy (ajprincess):

Area of a circle = \(\pi r^2\) r-radius

OpenStudy (anonymous):

The answers I can pick from are A.10 pi. B. 25pi. C. 50pi. or D. 100pi. And I did Pi r squared and it wasnt the answer I was looking for.

OpenStudy (ajprincess):

Area=\(\Large \pi r*r\)

OpenStudy (anonymous):

Its the same answer and I cant pick that answer...

OpenStudy (ajprincess):

what is the answer u r getting?

OpenStudy (anonymous):

78.5 is the answer I am getting but I cant use that answer

OpenStudy (anonymous):

25pi

OpenStudy (ajprincess):

well leave the \(\pi\) as it is and plug in 5 in place of r nd find the value

OpenStudy (anonymous):

\[A = \pi r squared\]

OpenStudy (anonymous):

thx 4 medal

OpenStudy (anonymous):

Welcome :)